OpenStudy (anonymous):

Find the prime factorization of 50.

OpenStudy (goformit100):



OpenStudy (anonymous):

what????

terenzreignz (terenzreignz):

50 = 2*25 25 = 5*5 50 = 2*5*5, 2 and 5 are both prime

OpenStudy (anonymous):

does it matter if they are prime?

OpenStudy (anonymous):

2*5*5 = 2*5^2 (3 prime factors, 2 distinct)

OpenStudy (anonymous):

\[2*5^2\] Answer. :-)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Make sure you also label the # of total factors and # of distinct factors

OpenStudy (anonymous):

Or it's technically only partly correct

OpenStudy (anonymous):

ok thanks for the help